OpenStudy (anonymous):

Will give you a medal for help on this. For the following pair of similar triangles, solve for x.

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry i cant help

OpenStudy (freckles):

so that means x+6 is similar to 39 and 2x-4 is similar to 24 so you solve: \[\frac{x+6}{24}=\frac{39}{2x-4}\]

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

So... first you level the equation our right? turn it into something that isn't a fraction?

OpenStudy (anonymous):

Like... \[(x+6) (2x-4) = 39 \times \]

OpenStudy (freckles):

(x+6)(2x-4)=39*24

OpenStudy (anonymous):

(x+4)(2x-4)=936 Whats the next step here??

OpenStudy (anonymous):

@freckles